• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Ran

    Log entry time 23:53:56 on January11,2005

    Entry number 139027

    keyword=shift summary

    17:00 asking MCC for Beam (done with all fixin the Hall)
    (needed to power cycle ROC4 and re-route N1-1L HV channel from 4.0L
    to 7.11U due to short-cut in cable.)

    17:20 - got Beam

    17:38 finished tilted carbon run with 250000 events

    18:40 moved to 10 microAmp beam

    18:41 starting to take DL2 data

    20:00 looking at Ctime , noticed we have a problem in the coinc time window.

    21:00 after consulting with Bob M it was suggested that we need to change our timming , we called MCC and asked for controled access.

    22:15 fix done .. now waiting for beam

    22:20 Beam is back 10 microAmp rastered

    24:00 still taking data